OpenStudy (anonymous):

Suppose you earn $12 each time you mow the lawn. Use function notation to describe the relationship between your total earnings E and the number of times m you mow the lawn.

OpenStudy (anonymous):

A)E(m) = m + 12 B)m = 12E C)E(m) = m – 12 D)E(m) = 12m answers

OpenStudy (sw050399):

Function notation simply means that you're going to have something like, E(m) where m is the variable you should see in the problem. Well, if you earn $12 per lawn mowing, m is the amount of times you mow the lawn, and E is the total earnings, then what operation will you be using?

OpenStudy (anonymous):

E(m)=12m ?

OpenStudy (sw050399):

Yep.